Crane sells T-shirts imprinted with high school names and logos. Last year the shirts sold for $26 each, and variable costs were $14.00 per shirt. At this cost structure, the breakeven point was 45,000 shirts. However, the company actually earned $21,600 in net income.

This year, the company is increasing its price to $28 per shirt. Variable costs per shirt will increase by Unresolved%, and fixed expenses will increase by $1,107,240. The tax rate will remain at 40%.Your answer is partially correct. Prepare a contribution format income statement for last year. (Round per unit computations
Answer #1

Last Year:

Contribution per unit = selling price per unit - variable costs per unit

Contribution per unit = $26 -$14 = $12

P/V Ratio = Contribution per unit / Selling price per unit *100

P/V Ratio = $12 / $26 *100 = 46.15%

Break Even Point Sales(Units) = Fixed Cost / Contribution per unit

45,000 = Fixed Cost / $12

Fixed Cost = 45,000 * $12

Fixed Cost = $540,000

Net Income = Operating Income ( 1 - 0.40)

$21,600 = Operating Income (1 - 0.40)

Operating Income = $21,600 / (1 - 0.40)

Operating Income = $21,600 / 0.60

Operating Income = $36,000

Operating Income + Fixed Cost = Contribution Margin

Contribution Margin = Operating Income + Fixed Cost

Contribution Margin = $36,000 +  $540,000

Contribution Margin = $576,000

Contribution Margin = Sales Qty * Contribution per unit

$576,000 / $12 = Sales Quantity

Sales Quantity = 48,000 units

Total Sales = 48,000 units * $26

Total Sales = $1,248,000

Total Variable cost = 48,000 units * $14

Total Variable cost = $672,000

Contribution Income Statement for the last year

Particulars Amount($) Per Unit

Sales $1,248,000 $26

Less: Variable Cost $672,000 $14

Contribution Margin $576,000 $12

Less: Fixed Cost $540,000

Operating Profit       $36,000

Less: Income Tax(40%) $14,400

Net Income $21,600